# Incremental Rotation Averaging Revisited

### 25 Dec 2023 v2

@@ Abstract
In order to further advance the accuracy and robustness of the incremental parameter estimation-based rotation averaging methods, in this paper, a new member of the Incremental Rotation Averaging (IRA) family is introduced, which is termed as IRAv4. As its most significant feature, a task-specific connected dominating set is extracted in IRAv4 to serve as a more reliable and accurate reference for rotation local-to-global alignment. This alignment reference is incrementally constructed, together with the absolute rotations of the vertices belong to it simultaneously estimated. Comprehensive evaluations are performed on the 1DSfM dataset, by which the effectiveness of both the reference construction method and the entire rotation averaging pipeline proposed in this paper is demonstrated.
@@
Global structure from motion, large-scale rotation averaging, task-specific connected dominating set.I Introduction
Image-based large-scale scene 3D reconstruction is a fundamental task in computer vision community, which has been widely investigated in recent years [1, 2, 3, 4]. As its core step, Structure from Motion (SfM) [5, 6, 7, 8, 9, 10, 11] aims to simultaneously recover camera pose and scene structure given pair-wise image feature matches. According to the camera pose initialization scheme, SfM methods could be roughly divided into incremental and global ones. The camera poses are sequentially estimated in incremental SfM via a iterative optimization pipeline [6, 8], but are simultaneously solved in global SfM by the motion averaging methodology [5, 9]. Compared with incremental SfM methods, global ones feature fewer optimization iterations and estimation parameters, which makes them with more advantage in theory and greater potential for application when dealing with increasingly larger reconstruction scene scales.
Motion averaging, which takes relative camera motions (rotations and translations) as input and produces absolute camera poses (orientations and locations), is the primary technique used for camera pose recovery in global SfM. Due to scale ambiguity of the relative translation estimated via essential matrix decomposition [12], motion averaging is conducted in the manner of first rotation and then translation averaging in most cases. As the former of the above two phases, rotation averaging [13] directly influences the effect of the subsequent translation averaging phase, and even all the remaining 3D reconstruction procedure, e.g. multi-view triangulation for scene recovery and global Bundle Adjustment (BA) for parameter optimization. Though increasing attention has been drawn in recent years [14, 15, 16, 17, 18, 19, 20], the problem of rotation averaging is still far from being solved due to the issues of large scale, imbalanced connectivity, and high-level noise of the Epipolar-geometry Graph (EG).
@@
However, in this paper we argue that the accuracy and robustness of the IRA series described above could be further advanced. Though the effectiveness of the cluster-based pipeline with task-independent CDS serving as global reference has been demonstrated in both IRAv3+ [24] and Jiang et al. [25], we believe that with a task-specific CDS extracted, more reliable global reference construction together with more accurate camera pose globalization would be further achieved. Based on the above analysis, a novel rotation averaging method termed as IRAv4 is proposed in this paper, which is built upon IRAv3+. The major difference between them lies in that instead of first extracting multiple CDSs and then estimating the absolute rotations in the CDSs’ local coordinate system by leveraging IRA, which is done in IRAv3+, the (task-specific) CDS is extracted by incrementally selecting the Next-Best Vertex (NBV) to maximize the supports from the currently extracted ones in the CDS, together with its absolute rotation (in the local coordinate system of CDS) simultaneously estimated. Our IRAv4 is evaluated on the 1DSfM dataset and compared with several other mainstream rotation averaging methods to demonstrate its effectiveness and advantages.
II Related Works
Except for the IRA series, there are mainly three kinds of large-scale rotation averaging methods, including the robust loss-based [26, 27, 28], the outlier filtering-based [29, 30, 31], and the deep learning-based [32, 33, 34] ones, which are briefly reviewed in the following.
II-A Robust Loss-Based Rotation Averaging
The robust loss-based methods [26, 27, 28] seek to design robust loss functions to achieve a robust estimation of the absolute camera rotations in the presence of relative rotation measurement outliers. Chatterjee et al. [26] developed a two-stage rotation averaging method. They firstly use the solution as an initialization to make the method robust to outliers. Then, an Iteratively Reweighted Least Squares (IRLS) approach based on the loss function is followed to achieve an efficient and accurate absolute rotation estimation. Shi et al. [27] proposed a novel robust optimization framework, Message Passing Least Squares (MPLS), as an alternative paradigm to IRLS for rotation averaging, and the advantages of MPLS over IRLS are theoretically analyzed. After that, based on the cyclic consistency constraint, Shi et al. [28] presented a novel quadratic programming framework, Detection and Estimation of Structural Consistency (DESC), which is less dependent to initialization and more simple in optimization. Though the above methods are robust to the relative rotation outliers to some extent, the optimization procedures in them still heavily depend on the initialization and are usually inefficient because of the relatively complicated problem formulations.
II-B Outlier Filtering-Based Rotation Averaging
Compared with the robust loss-based methods, the outlier filtering-based ones [29, 30, 31], are more theoretically straightforward. Cui et al. [29] proposed a multiple Orthogonal Maximum Spanning Trees (OMSTs)-based rotation averaging method. Multiple OMSTs are selected from the input EG to obtain a set of densely connected and more accurate relative rotation measurements for robust rotation averaging [26]. Gao et al. [30] presented a Hierarchical RANSAC-based Rotation Averaging (HRRA) method. The absolute rotations are estimated based on the thought of RANSAC with random spanning tree of the input EG serving as the minimal set for model estimation. In addition, graph clustering is employed to downsize the spanning trees as the RANSAC-based parameter estimation is sensitive to the size of the minimal set. Lee et al. [31] introduced the algorithm of HARA, which incrementally adds constraints based on the size of the supporting camera triplets to construct EG’s spanning tree and remove outliers in relative rotation measurements, thus achieving a hierarchical and robust initialization of camera absolute rotation. Though effective outlier filtering and accurate rotation estimation could be achieved in most cases, we find that the performance of the above methods exhibit some instability when dealing with the EGs with relatively high noise level.
II-C Deep Learning-Based Rotation Averaging
Recently, several deep learning-based methods have been proposed dealing with the rotation averaging problem by leveraging graph neural networks [32, 33, 34]. Purkait et al. [32] proposed NeuRoRA, a two-stage graph neural networks for rotation averaging. The first stage is a view-graph cleaning network to detect relative rotation measurement outliers. The second stage is a fine-tuning network to refine the absolute orientations. Yang et al. [33] presented Multi-Source Propagator (MSP), which takes appearance information as input and introduces a differentiable Shortest Path Tree (SPT) method to achieve a robust initialization result. The initialization is further improved through non-learnable iterative edge reweighting. Li et al. [34] proposed the method of Rotation Averaging Graph Optimizer (RAGO), a learnable iterative graph optimizer minimizing a gauge-invariant cost function with an edge rectification strategy to mitigate the effect of inaccurate measurements. However, these deep learning-based methods depend heavily on the training data, which makes their generalization and scalability poor in novel situations.
III Brief Description on the previous IRA Series
@@
IV The Proposed IRAv4 Method
@@
IV-A Task-Specific Alignment Reference Construction
@@
the above optimization problem and the other ones in the rest of this paper are solved by the Ceres Solver library, and the triplet with the largest selection reward defined in the following, together with the optimized rotations are served as the initial seed construction:
@@
where is the initialization of the vertex . Then, the absolute rotation of is locally optimized by:
@@
IV-B Other Key Steps Stemming from the IRAv3+
@@
Data | |||||||
ALM | |||||||
ELS | |||||||
GDM | |||||||
MDR | |||||||
MND | |||||||
NYC | |||||||
PDP | |||||||
PIC | |||||||
ROF | |||||||
TOL | |||||||
TFG | |||||||
USQ | |||||||
VNC | |||||||
@@
Data |
Robust loss-based | Outlier filtering-based | Deep learning-based | Previous IRA series | IRAv4 | |||||||||
14 | ||||||||||||||
IRLS- |
MPLS |
DESC |
OMSTs |
HRRA |
HARA |
NeuRoRA |
MSP |
RAGO |
IRA |
IRA++ |
IRAv3 |
IRAv3+ |
[26] |
[27] |
[28] |
[29] |
[30] |
[31] |
[32] |
[33] |
[34] |
[21] |
[22] |
[23] |
[24] |
ALM |
0.73° | |||||||||||||
ELS |
GDM |
MDR |
MND |
NYC |
PDP |
PIC |
ROF |
TOL |
TFG |
USQ |
VNC |
YKM |
@@
V Experimental Evaluation
In this section, the dataset and metric for evaluation are firstly described, and then the proposed IRAv4 method is evaluated in terms of absolute rotation estimation and alginment reference construction for its effectiveness demonstration.
V-A Dataset and Metric for Evaluation
The datasets for method evaluation employed in this paper are the commonly-used 1DSfM dataset [37], which contains groups of test data stemmed from the Internet. The metadata of the dataset is listed in Table I, where the absolute rotation ground truths of each test data are provided by the SfM results of Bundler [38]. The accuracy evaluation metric used in this paper is the same as that in the previous IRA series [21, 22, 23, 24]. Specifically, a best relative rotation between the absolute rotation estimations and ground truths is estimated and used to align these two rotation sets. And then the median alignment error in degree is used as the evaluation metric.
V-B Comparison in Absolute Rotation Estimation
Except for the IRA series, including IRA [21], IRA++ [22], IRAv3 [23], IRAv3+ [24], and IRAv4 proposed in this paper, which have been described in Sec. III and Sec. IV, there are some other mainstream rotation averaging methods involved for absolute rotation estimation comparison, including robust loss-based methods, IRLS- [26], MPLS [27], DESC [28], outlier filtering-based methods, OMSTs [29], HRRA [30], and HARA [31], and deep learning-based methods, NeuRoRA [32], MSP [33], and RAGO [34]. The comparative results are illustrated in Table IV-B, and to deliver a comprehensive comparison, the rankings of the comparative methods on each test data are averaged and shown in the last row of the table. It could be observed that IRAv4 achieves overall best performance among all the comparative methods, which demonstrates its advantages in absolute rotation estimation.
In addition, to give a more comprehensive comparison on different rotation averaging methods, the absolute rotations obtained by them are fed into a typical global SfM pipeline for sparse scene reconstruction, including translation averaging (BATA [39]) and multi-view triangulation [6] for the initialization of camera locations and scene points, and global bundle adjustment [40] for camera pose and scene structure optimization. The sparse reconstruction results of GDM with symmetrical structure by IRLS-, HARA, NeuRoRA, and IRAv4-based global SfM are shown in Fig. 2. It could be observed that among the global SfM variants, only the one based on our IRAv4 could successfully recover the correct scene structure of GDM, which validates its robustness to scenes with ambiguous structures and EG with high noise.

### 05 Jan 2024 v3

@@ Abstract
In order to further advance the accuracy and robustness of the incremental parameter estimation-based rotation averaging methods, in this paper, a new member of the Incremental Rotation Averaging (IRA) family is introduced, which is termed as IRAv4. As the most significant feature of the IRAv4, a task-specific connected dominating set is extracted to serve as a more reliable and accurate reference for rotation global alignment. In addition, to further address the limitations of the existing rotation averaging benchmark of relying on the slightly outdated Bundler camera calibration results as ground truths and focusing solely on rotation estimation accuracy, this paper presents a new COLMAP-based rotation averaging benchmark that incorporates a cross check between COLMAP and Bundler, and employ the accuracy of both rotation and downstream location estimation as evaluation metrics, which is desired to provide a more reliable and comprehensive evaluation tool for the rotation averaging research. Comprehensive comparisons between the proposed IRAv4 and other mainstream rotation averaging methods on this new benchmark demonstrate the effectiveness of our proposed approach.
@@
Large-scale rotation averaging, task-specific connected dominating set, rotation averaging benchmarking.I Introduction
Image-based large-scale scene 3D reconstruction is a fundamental task in computer vision community, which has been widely investigated in recent years [1, 2, 3, 4]. As its core step, Structure from Motion (SfM) [5, 6, 7, 8, 9] aims to simultaneously recover camera pose and scene structure given pair-wise image feature matches. According to the camera pose initialization scheme, SfM methods could be roughly divided into incremental and global ones. The camera poses are sequentially estimated in incremental SfM via a iterative optimization pipeline [6, 8], but are simultaneously solved in global SfM by the motion averaging methodology [5, 9]. Compared with incremental SfM methods, global ones feature fewer optimization iterations and estimation parameters, which makes them with more advantage in theory and greater potential for application when dealing with increasingly larger reconstruction scene scales.
Motion averaging, which takes relative camera motions (rotations and translations) as input and produces absolute camera poses (orientations and locations), is the primary technique used for camera pose recovery in global SfM. Due to scale ambiguity of the relative translation estimated via essential matrix decomposition [10], motion averaging is conducted in the manner of first rotation and then translation averaging in most cases. As the former of the above two phases, rotation averaging [11] directly influences the effect of the subsequent translation averaging phase, and even all the remaining 3D reconstruction procedure, e.g. multi-view triangulation for scene recovery and global Bundle Adjustment (BA) for parameter optimization. Though increasing attention has been drawn in recent years [12, 13, 14, 15, 16, 17, 18, 19, 20], the problem of rotation averaging is still far from being solved due to the large scale, imbalanced connectivity, and high-level noise in the Epipolar-geometry Graph (EG).
@@
However, in this paper we argue that the accuracy and robustness of the IRA series described above could be advanced one step further. Though the effectiveness of the cluster-based pipeline with task-independent CDS serving as global reference has been demonstrated in both IRAv3+ [24] and Jiang et al. [25], we believe that with a task-specific CDS extracted, more reliable global reference construction together with more accurate camera pose globalization would be further achieved. Based on the above analysis, a novel rotation averaging method termed as IRAv4 is proposed in this paper, which is built upon IRAv3+. The major difference between them lies in that instead of first extracting multiple CDSs and then estimating the absolute rotations in the CDSs’ local coordinate system by leveraging IRA, which is done in IRAv3+, the (task-specific) CDS is extracted by incrementally selecting the Next-Best Vertex (NBV) to maximize the supports from the currently extracted ones in the CDS, together with its absolute rotation (in the local coordinate system of CDS) simultaneously estimated.
Moreover, we find two issues that need to be addressed when evaluating the performance of most existing rotation averaging methods [26, 27, 28, 29, 16, 21, 30, 19, 31, 22, 23, 24]. Firstly, the evaluation is mostly performed based on the 1DSfM [32] dataset with the camera calibration results of the slightly outdated Bundler [33] serving as ground truth. And secondly, the evaluation mostly focuses on the camera rotation estimation accuracy only while ignoring the influence of the estimated rotations to the downstream task, i.e. translation averaging. To deal with these, a new COLMAP-based [6] rotation averaging benchmark is rebuilt upon the 1DSfM dataset in this paper, where the EG of each test data in the 1DSfM dataset111https://www.cs.cornell.edu/projects/1dsfm/ is regenerated by leveraging COLMAP222https://demuc.de/colmap/ and OpenCV333https://opencv.org/ libraries. To provide a more reliable ground-truth source, only the camera poses pass the cross check between COLMAP and Bundler are employed in this new benchmark. In order to additionally evaluate the effectiveness of the estimated rotations for the downstream translation averaging task, in our new benchmark, the rotations estimated by different rotation averaging methods, together with the relative translations in the regenerated EG, are feed into a well-established translation averaging method, BATA [34], to obtain absolute camera locations for location estimation accuracy evaluation. Based on the new benchmark, comprehensive evaluation is performed on the proposed IRAv4 and several currently mainstream rotation averaging methods, including the existing IRA series [21, 22, 23, 24], and some other methods [35, 26, 27, 30, 31, 28]. Among all these methods for performance evaluation, the proposed IRAv4 in this paper achieves state of the art in both rotation estimation and the downstream location estimation on the new benchmark, by which its effectiveness is demonstrated.
The main contributions of this paper are threefold:
1) A novel cluster-based rotation averaging pipeline is proposed, where a task-specific CDS is extracted to serve as the global reference for rotation local-to-global alignment. 2) A new rotation averaging benchmark based on the 1DSfM dataset and the COLMAP library is presented, where both the ground truth source and the performance evaluation metrics are redefined for a more reliable and comprehensive evaluation. 3) A comprehensive evaluation is conducted between the proposed IRAv4 and several mainstream rotation averaging methods on the new benchmark, where state-of-the-art performance of IRAv4 is achieved to demonstrate its effectiveness.
II Brief Description on the IRA Series
@@
III The Proposed IRAv4 Method
@@
III-A Task-Specific CDS-Based Global Reference Construction
@@
the above optimization problem and the other ones in the rest of this paper are solved by the Ceres Solver library444http://www.ceres-solver.org/, and the triplet with the largest selection reward defined in the following, together with the optimized rotations are served as the initial seed construction:
@@
where is the rotation initialization of the vertex . Then, the absolute rotation of is locally optimized by:
@@
III-B Other Key Steps Stemming from the IRAv3+
@@
IV New Rotation Averaging Benchmark Creation
It could be observed that most of the existing rotation averaging methods [26, 27, 28, 29, 16, 21, 30, 19, 31, 22, 23, 24] perform performance evaluation based on the 1DSfM [32] dataset with the slightly outdated Bundler’s [33] camera calibration results serving as the ground-truth source of camera rotations, which has been presented for more than ten years. In order to give a more comprehensive evaluation on the existing mainstream rotation averaging methods, a new 1DSfM-derived benchmark based on COLMAP [6] is rebuilt here. COLMAP is a general-purpose, end-to-end image-based 3D reconstruction pipeline with a graphical and command-line interface, which is proposed more recently and updated more frequently. In recent years, COLMAP has become the most widely recognized and commonly used open-source SfM toolbox with various downstream applications, such as novel view synthesis [38, 39] and neural surface reconstruction [40, 41]. As a result, we believe that it is time for COLMAP to take the place of Bundler for rotation averaging benchmarking. The new benchmark creation process mainly contains three parts: epipolar-geometry graph regeneration, ground-truth data acquisition, and evaluation metrics definition, which are thoroughly described in the following, respectively.
IV-A Epipolar-Geometry Graph Regeneration
In order to present a new rotation averaging benchmark by making full use of COLMAP, the Epipolar-geometry Graph (EG) of each test data in the 1DSfM dataset is regenerated at first. It should be noted that only the images in the Maximum Connected Component (MCC) are considered during the EG regeneration procedure here for the efficiency consideration, and this image connectivity information is provided by the file contained in the original 1DSfM dataset. Then, the images to be processed are fed into the COLMAP toolbox for local feature extraction and matching with default parameters. Given the pair-wise feature matches produced by COLMAP, OpenCV library is employed for essential matrix estimation and decomposition to produce relative motion (rotation and translation), where the feature-level point to epipolar line distance threshold and the desirable level of estimation confidence are set to and respectively during the RANdom SAmple Consensus (RANSAC) [42] procedure of the essential matrix estimation [10] process. After that, the image pair with more than pairs of inlier feature matches for RANSAC-based essential matrix estimation is added as a graph edge to the regenerated EG. It should be further noted that the camera intrinsic parameters are required during essential matrix estimation, which are provided by the file contained in the original 1DSfM dataset here.
Data | |||||||||||||
ALM | |||||||||||||
ELS | |||||||||||||
GDM | |||||||||||||
MDR | |||||||||||||
MND | |||||||||||||
NYC | |||||||||||||
PDP | |||||||||||||
PIC | |||||||||||||
ROF | |||||||||||||
TOL | |||||||||||||
TFG | |||||||||||||
USQ | |||||||||||||
VNC | |||||||||||||
@@ IV-A Epipolar-Geometry Graph Regeneration
IV-B Ground-Truth Data Acquisition
After EG regeneration, COLMAP is employed again for the ground-truth data acquisition of camera absolute rotations (and locations) for rotation averaging evaluation. Specifically, for each test data of the 1DSfM dataset, given the image matching result of COLMAP itself obtained in the last subsection, the SfM procedure integrated in COLMAP is conducted with default parameters, and the camera calibration result is denoted as , where is the set of cameras successfully registered by COLMAP. In addition, the camera calibration result of Bundler, provided by the file contained in the original 1DSfM dataset, is denoted as . It should be noted that has been aligned to the GPS coordinate system and is with real physical scale, while suffers from the issue of scale ambiguity due to the metric reconstruction nature of SfM. As a result, in order to use the camera calibration result of COLMAP as ground-truth camera poses for rotation (and translation) averaging evaluation, it should be aligned to the coordinate system of Bundler for scale recovery. The alignment is achieved by estimating a similarity transformation [43] based on RANSAC with the co-calibrated cameras by COLMAP and Bundler, , where the locational distance threshold and the desirable estimation confidence are set to and , respectively. Then, based on the estimated transformation , could be aligned to by:
Before performing rotation averaging method benchmarking, it should be further noticed that the basic assumption of involving the camera calibration result of the incremental SfM methods (Bundler or COLMAP) to serve as ground-truth source for motion (rotation and translation) averaging evaluation lies in that the camera orientation and localization results of incremental SfM with iterative RANSAC-based outlier filtering and BA-based parameter optimization [44] are acknowledged to be with much higher accuracy than those of motion averaging before performing the final global BA. Nonetheless, the incremental SfM-based calibration result inevitably contains estimation errors and is still not the real ground-truth camera poses. To make the evaluation with more reliability, inspired by the work of long-term visual localization benchmarking [45], the camera poses calibrated by COLMAP that pass Bundler-based cross check with certain pose accuracy level are used as the final camera pose ground-truth data. Its basic idea lies in that the camera poses calibrated by both COLMAP and Bundler with enough closeness in the -dimensional camera pose space are used for the following rotation averaging evaluation. Specifically, given the rotational () and locational () distance thresholds with different accuracy levels of camera poses: high-accuracy , medium-accuracy , and coarse-accuracy , the camera subsets for rotation averaging evaluation with respect to the different distance thresholds are obtained by:
The sizes of the camera subsets together with some other meta-data about the new rotation averaging benchmark are listed in Table I, and in order to make a trade-off between ground-truth data magnitude and reliability, the camera subset with medium-accuracy, i.e. , are finally employed for rotation averaging benchmarking. The COLMAP-to-Bundler alignment median errors of rotation and location in Table I are defined as:
where returns the median value of the set . It could be further observed from the table that: 1) in most cases, except for MDR, TOL, and YKM, more cameras could be registered by COLMAP than Bundler; 2) the disparity in camera calibration accuracy between COLMAP and Bundler is insignificant as both and are relatively small; 3) Different test data in the new benchmark is with different degree of conformity in accuracy () between COLMAP and Bundler, take for example, co-calibrated cameras pass the median-level-accuracy cross check for MND, while for TFG, that value decreases to only .
Data | IRLS-GM | IRLS- | MPLS | DESC | HARA | NeuRoRA | IRA | IRA++ | IRAv3 | IRAv3+ | IRAv4 | MPLS | IRAv4 |
[35] | [26] | [27] | [31] | [30] | [28] | [21] | [22] | [23] | [24] | w/ BA | w/ BA | ||
ALM | |||||||||||||
ELS | |||||||||||||
GDM | |||||||||||||
MDR | |||||||||||||
MND | |||||||||||||
NYC | |||||||||||||
PDP | |||||||||||||
PIC | |||||||||||||
ROF | |||||||||||||
TOL | |||||||||||||
TFG | |||||||||||||
USQ | |||||||||||||
VNC | |||||||||||||
YKM | |||||||||||||
@@ IV-B Ground-Truth Data Acquisition
IV-C Evaluation Metrics Definition
For a particular rotation averaging method (what could be is described in the next section), it is firstly performed on the relative rotations from the regenerated EG in this paper to obtain the absolute camera rotations. Then, based on the rotation averaging result and the relative translation measurements, translation averaging is performed with the mainstream method, BATA555https://bbzh.github.io/ [34], to get the absolute camera locations, and the -based motion averaging result is denoted as . Then, the rotation and similarity transformation for aligning to the coordinate system of are estimated by the methods of voting-based single rotation averaging [29] and RANSAC-based similarity estimation, respectively. The pose alignment is conducted in the same way of Eq. 13, and the aligned poses are denoted as . Finally, the median rotational and locational errors on the camera subset of are served as the evaluation metrics of the new rotation averaging benchmark, which are defined as:
V Rotation Averaging Method Benchmarking
In this section, several mainstream rotation averaging methods for comparison are first introduced, and then rotation averaging method benchmarking in terms of rotation and location estimation is performed based on the presented new rotation averaging benchmark.
Data | IRLS-GM | IRLS- | MPLS | DESC | HARA | NeuRoRA | IRA | IRA++ | IRAv3 | IRAv3+ | IRAv4 | MPLS | IRAv4 |
[35] | [26] | [27] | [31] | [30] | [28] | [21] | [22] | [23] | [24] | w/ BA | w/ BA | ||
ALM | |||||||||||||
ELS | |||||||||||||
GDM | |||||||||||||
MDR | |||||||||||||
MND | |||||||||||||
NYC | |||||||||||||
PDP | |||||||||||||
PIC | |||||||||||||
ROF | |||||||||||||
TOL | |||||||||||||
TFG | |||||||||||||
USQ | |||||||||||||
VNC | |||||||||||||
YKM | |||||||||||||
Rank |
V-A Rotation Averaging Methods for Comparison
Except for the IRA series (IRA [21], IRA++ [22], IRAv3 [23], IRAv3+ [24], and IRAv4 proposed in this paper) which have been described in Sec. II and Sec. III, there are some other mainstream rotation averaging methods involved in this section for comparison, including IRLS-GM666https://ee.iisc.ac.in/cvlab/research/rotaveraging/ [35], IRLS-6 [26], MPLS777https://github.com/yunpeng-shi/MPLS [27], DESC888https://github.com/ColeWyeth/DESC [31], HARA999https://github.com/sunghoon031/HARA [30], and NeuRoRA101010https://github.com/pulak09/NeuRoRA [28]. These methods are briefly introduced in the following for better understanding.
IRLS-GM [35] works on the Lie group structure of rotations and solves the problem of large-scale robust rotation averaging in a two-stage manner. -based optimization is first carried out to achieve an efficient, scalable, and robust solution. Based-on the -based initialization, Iterative Reweighted Least Squares (IRLS)-based optimization with the Geman-McClure (GM) loss function is performed to further improve the accuracy of the solution.
IRLS- [26] is the upgrade version of IRLS-GM, which also contains two steps. IRLS-GM and IRLS- share the same step one. For the second step of IRLS-, a fairly exhaustive evaluation of the performance of a variety of robust loss functions that are popular in the statistical literature is provided, which leads to a recommendation of the loss function that provides the best empirical performance.
MPLS [27] is a framework that could be served as alternative paradigm to IRLS for rotation averaging under high corruption and noise. It uses the theoretically guaranteed cycle-edge message passing algorithm for estimating the underlying corruption levels. These estimates are then used for learning better weights for the weighted least squares problem. The limitations of IRLS and the advantages of MPLS for solving the rotation averaging are also discussed and analyzed in the original paper of MPLS.
DESC [31] is a novel quadratic programming formulation for estimating the corruption levels in rotation averaging, which are used for solving this problem. Its objective function exploits the cycle consistency of the Lie group and has the following advantages: 1) It can tolerate corruption as high as the information-theoretic bound; 2) It does not require a good initialization for the estimates of group elements; 3) It has a simple interpretation; And 4) under some mild conditions its global minimum exactly recovers the corruption levels.
HARA [30] presents a hierarchical initialization scheme that constructs a spanning tree of a rotation graph by propagating most reliable constraints first and less reliable ones later. In HARA, the hierarchy of reliability based on the number of consistent triplet constraints, as well as their level of consistency, is established. That is, a constraint is considered to be more reliable if it is strongly supported by many other constraints and less reliable if it has weaker or fewer supports. In addition, the number of valid 2D-2D correspondences could also be optionally incorporated into the hierarchy.
NeuRoRA [28] is built to learn patterns from the data and predict/regress the model parameters from the noisy relative rotation measurements. NeuRoRA is a two-step approach: In the first step, a graph-based network, CleanNet, is utilized to clean the EG by removing outliers and rectifying noisy measurements; And in the second step, an initialization from the cleaned EG, instantiated from a shortest path tree, is then further fine-tuned using a separate graph-based network, FineNet, to produce the final rotation averaging results.
V-B Rotation Estimation Accuracy Benchmarking
For the methods of IRLS-GM, IRLS-, MPLS, DESC, and HARA, we locally run the official source codes (with default parameters) released by the authors on the new 1DSfM-derived rotation averaging benchmark to produce the absolute rotation estimation results of them. For the method of NeuRoRA, follow the authors’ instruction, the CleanNet and FineNet of NeuRoRA are both firstly pretrained on synthetic data and then finetuned and tested on the real-world data (with default parameters) in round-robin fashion (leave one out). And for the methods of existing IRA series (IRA, IRA++, IRAv3, and IRAv3+) and the proposed IRAv4, locally implemented source codes are employed for absolute rotation estimation. The absolute rotation estimation results of all the methods for comparison on the new benchmark are shown in Table II. The top-four results on each test data among all the comparative methods are highlighted, and in order to perform a comprehensive comparison, the rankings of all the comparative methods on each test data are averaged and shown in the last row of the table, and the top-four methods are highlighted as well. It could be seen from the table that: 1) Among all the rotation averaging methods for comparison, the proposed IRAv4 in this paper achieves overall the best performance in rotation estimation accuracy on the new 1DSfM-derived rotation averaging benchmark; 2) The deep learning-based rotation averaging method NeuRoRA is rather data/domain-sensitive and performs relatively poor in rotation estimation accuracy on the new benchmark, especially for the large-scale (PIC, ROF, and TFG) and high-noise (GDM, MDR, and USQ) test data; And 3) Compared with the results of IRA++ and IRAv3+, better performance is achieved by IRAv3 and IRAv4 in rotation estimation accuracy on the new benchmark, which demonstrates the effectiveness of the task-specific epipolar-geometry graph clustering and alignment reference construction proposed in IRAv3 and IRAv4, respectively.
In addition, in order to further demonstrate the effectiveness of the proposed IRAv4, comparative experiments in rotation estimation accuracy on the original 1DSfM dataset are also conducted, and the results are shown in Table III. It could be observed that compared with the state-of-the-art method on the dataset, RAGO [19], and the existing IRA series, IRAv4 is able to achieve overall the best performance as well. And the reasons for the inconsistency of different methods in rotation estimation accuracy on the two benchmarks (cf. Table II and Table III) mainly lie in the differences in ground-truth rotation source (different number of cameras involved in evaluation and different ways in obtaining their ground-truth rotations).
V-C Location Estimation Accuracy Benchmarking
Given the absolute rotation estimation results of all the rotation averaging methods for comparison on the new 1DSfM-derived benchmark (and the relative translations in the regenerated EG, which is mandatory for location estimation), we locally run the official source code of BATA (with default parameters) released by the authors to produce the absolute location estimation results of these methods. The absolute location estimation results of all the methods for comparison on the new benchmark are shown in Table IV. And similar to that in Table II, the top-four methods on each test data and in terms of overall rankings are highlighted. By observing the above two tables, similar conclusions, i.e. the three items discussed in the last sub-section, could be drawn. It could be further observed that though with moderately strong correlation, among different rotation averaging methods, the advantage in rotation estimation accuracy would not be strictly maintained during location estimation, e.g. the third and fourth-best methods in rotation estimation accuracy on the new benchmark are IRAv3 and IRLS-, while those in location estimation accuracy are IRA and IRLS-GM.
V-D Incorporation of the Global Bundle Adjustment
In order to demonstrate the performance of different rotation averaging methods coupled with some certain translation averaging method (BATA) in providing accurate enough camera pose initializations for the final BA, multi-view triangulation-based scene computation and global BA-based pose optimization are performed. Specifically, the pair-wise feature matches those pass through the RANSAC-based essential matrix estimation are linked to feature tracks with the union find algorithm implemented in OpenMVG111111https://github.com/openMVG/openMVG. Then, a RANSAC-based multi-view triangulation algorithm implemented in COLMAP is performed with the reprojection error and viewing angle thresholds, and the desirable level of estimation confidence setting to , , and to compute the initial values of the feature tracks’ spatial coordinates. During the multi-view triangulation, the intrinsic parameters provided by the original 1DSfM dataset, and the camera poses estimated by different rotation averaging methods and BATA are used to formulate camera projection matrices. Finally, given the initial values of camera parameters and scene points, global BA is performed to further improve the camera pose accuracy, which is conducted by leveraging Ceres Solver with Huber loss. The rotation and location estimation accuracy of the optimized camera poses of the top-2 methods in both rotation and location estimation accuracy before global BA (IRAv4 and MPLS) on the new benchmark are shown in the last two columns of Table II and Table IV, respectively. Note that the results of them are re-ranked independently in the tables to maintain comparative fairness. It could be observed from the two tables that with the global BA, the camera pose estimation accuracy of both IRAv4 and MPLS is largely improved, and IRAv4 achieves slightly better performance in location estimation accuracy than MPLS, which demonstrates that accurate enough camera pose initializations cloud be provided by both IRAv4 and MPLS coupling BATA, and the advantage of IRAv4 over MPLS.
Furthermore, to provide a more intuitive comparison in robustness of all the rotation averaging methods for comparison, the sparse scene reconstruction results after global BA of them on the test data of GDM of the new rotation averaging benchmark are shown in Fig. 2. As shown in the upper-left of this figure, there are two buildings with left-right symmetry located in the scene of GDM, which significantly increases the difficulty of camera pose estimation and the subsequent scene structure recovery. It could be observed from the figure that among all rotation averaging methods involved in benchmarking, only MPLS, DESC, and the IRA series succeed to recovery the correct scene structure of GDM, which demonstrate their robustness.
VI conclusion
In this paper, to benefit from the global reference for local-to-global rotation alignment more, an incremental parameter estimation-based task-specific connected dominating set extraction method is proposed, by which the accuracy and robustness of the Incremental Rotation Averaging (IRA) series are further advanced. Furthermore, a new 1DSfM-derived rotation averaging benchmark, together with new ground-truth data acquisition approach, and new evaluation tasks and metrics, based on the camera calibration results of COLMAP is presented, in order to give a more comprehensive and reasonable comparison among different rotation averaging methods. Based on the new benchmark, the proposed method in this paper, IRAv4, achieves overall the best performance in both rotation and downstream location estimation, which demonstrates its effectiveness in terms of accuracy and robustness when dealing with large-scale and high-noise rotation averaging problems.